# NEMO-4-PAYPAL: Leveraging NVIDIA’s Nemo Framework for empowering PayPal’s Commerce Agent

### v2

@@ Abstract
Figure 1: PayPal is transforming from a payments company to a commerce platform with, our commerce Paypal Agent, which leverages NVIDIA’s Nemo.
@@ 1 Introduction
The emergence of agentic artificial intelligence represents a paradigm shift in how consumers interact with e-commerce platforms purcarea2025commerce; kodali2025agentic, promising to transform the traditional shopping experience into an intelligent, conversational journey. PayPal’s Commerce Agent embodies this vision by enabling autonomous shopping assistance across the entire commerce lifecycle, from pre-purchase discovery through post-purchase support. This multi-agent system leverages advanced language models to provide personalized shopping experiences, positioning itself at the forefront of the agentic commerce revolution.
The complexity of modern e-commerce interactions demands sophisticated AI systems capable of understanding nuanced customer intent, navigating vast product catalogs, and orchestrating seamless transactions. Traditional rule-based systems and simple chatbots fail to meet these requirements, necessitating the development of intelligent agents powered by large language models (LLMs) goldrevolutionizing. However, the deployment of such systems in production environments presents significant challenges in terms of latency, cost, and performance optimization, particularly when handling the scale and diversity of PayPal’s global commerce ecosystem.
@@
We collaborated with NVIDIA to solve these performance challenges by leveraging the NeMo Framework NVIDIA_NeMo_Framework and Nemotron model family with specialized, fine-tuned small language models (SLMs) optimized for commerce search tasks.
@@ 2.1 E-Commerce Information Retrieval
Traditional e-commerce information retrieval systems manning_introduction_2008 have predominantly relied on structured data representations, operating over well-defined product catalogs, taxonomies, and attribute-based schemas to enable search and recommendation functionalities. The conventional approach involves a systematic pipeline of converting unstructured textual data—such as product descriptions, user reviews, and query inputs—into structured formats through information extraction techniques, followed by conducting search operations over these normalized data structures. This methodology leverages relational databases, inverted indices, and rule-based matching algorithms to achieve precise retrieval based on explicit attribute matching and categorical filtering. While this structured approach provides deterministic results and enables efficient query processing through established indexing mechanisms, it faces inherent limitations in handling semantic nuances, contextual understanding, and the growing complexity of natural language queries that characterize modern e-commerce interactions wang2024rethinking.
Recent advances in large language models (LLMs) have demonstrated strong capabilities in natural language understanding and generation, showing new opportunities for bridging the gap between unstructured user queries and structured product representations gao2023precise. These models can effectively extract user intent and product attributes from conversational queries, and reformulate ambiguous or underspecified searches into precise, structured formats that align with product catalog schemas, thereby enhancing retrieval effectiveness in e-commerce systems zhang2025leser; jagerman2023query.
@@ 2.2 LLM Agent in Search and Recommendation
LLM agents huang2024understanding are transforming commerce search and recommendation systems by enabling more natural, context-aware interactions between users and e-commerce platforms. Unlike traditional keyword-based search systems that rely on exact matches and predefined filters, LLM agents leverage large language models to understand user intent across multiple dimensions, including implicit preferences, contextual constraints, and evolving needs throughout a shopping session. These agents perform sophisticated query understanding by extracting attributes (e.g., brand, color, price range, size), expanding queries with related terms and synonyms, and reformulating ambiguous requests into structured search parameters. Beyond search, LLM agents orchestrate multi-step recommendation workflows by reasoning over product catalogs, user purchase history, and real-time session data to generate personalized suggestions. They handle complex conversational scenarios such as comparison requests (”show me Nike shoes similar to what I bought last month but in blue”), constraint satisfaction (”find running shoes under $\mathdollar 100$ with good arch support”), and cross-category exploration (”I need shoes for hiking, what else should I get for my trip?”). By integrating retrieval systems with ranking models, LLM agents ground their responses in actual product inventory while maintaining conversational fluency, ultimately reducing search friction, increasing discovery, and improving conversion rates in modern commerce platforms.
@@ 2.3 NVIDIA’s NeMo Framework
NVIDIA NeMo Framework NVIDIA_NeMo_Framework is a complete AI platform that helps researchers and developers work with Large Language Models (LLMs), multimodal systems, and speech AI applications. The framework supports the entire process of building AI models, from preparing data and training models to putting them into production, and can run on different computing environments including local servers, data centers, or cloud services. NeMo includes tools for data preparation, efficient training methods like supervised fine-tuning and LoRA (a technique that reduces training time and resources), and multiple options for deploying trained models.
@@
Figure 3: NeMo Framework nvidia_nemo_framework_2024 NVIDIA NeMo is an end-to-end platform for developing custom generative AI, anywhere. It includes tools for training, fine-tuning, retrieval-augmented generation, guardrailing, data curation, and pretrained models. It has offerings across the tech stack, from frameworks to high-level API endpoints.
@@ 3.1 LLM-Powered Commerce Search &\& Recommendation
To leverage the powerful text generation and understanding capabilities of large language models (LLMs), we adopt the Hypothetical Document Embeddings (HyDE) gao2023precise approach for product retrieval. When a user prompt is given, the LLM first extracts (attribute, value) pairs from the natural language query and reformulates it into a structured format of (product category, attributes, values). The LLM then generates hypothetical product information that captures the user’s intent and desired product characteristics. This hypothetical product description, while potentially containing idealized or imprecise details, effectively encodes the relevance patterns of what the user is seeking. The structured query and hypothetical product information are then sent to our search engine, which retrieves matching products from our product catalog by identifying similar real products in the embedding space, effectively grounding the generated hypothetical description to actual inventory while filtering out any inaccurate details through the encoder’s dense representation.
@@ 3.2 Commerce Agent System Architecture
To further our agentic commerce initiatives, we developed a comprehensive agentic system to support conversational AI experiences for users through both web and mobile platforms. Our architecture employs a Java-based backend API server, utilizing the LangChain mavroudis2024langchain agent orchestration framework to develop intelligent products that leverage multiple language models and coordinate internal knowledge systems, domain services, and existing machine learning solutions. Our system implements a hybrid cloud solution that integrates PayPal’s internal technology stack with cloud platform services to ensure scalability, reliability, and performance for production deployment. (see Figure 4)
@@ 4.2 Evaluation Methodology
We finetuned and evaluated multiple model variants with different hyperparameter configurations. Our evaluation employs LLM-as-a-Judge gu2024survey, using pairwise and rubric-based comparative evaluations for qualitative analysis, to assess model performance.
@@ 4.3 Supervised Fine Tuning (SFT)
We leveraged NVIDIA’s NeMo SDK as the foundation for fine-tuning a family of small language models (SLMs), focusing on the llama3.1-nemotron-nano-8B-v1 architecture. We conducted comprehensive experiments training 20 LoRA-based model variants while systematically sweeping hyperparameter spaces including learning rates, optimizers (Adam, AdamW), cosine annealing schedules, and LoRA ranks. The NeMo Framework SDK significantly enhanced developer productivity by providing turnkey multi-GPU and multi-dimensional parallelism capabilities, reducing time-to-train through well-maintained training recipes. NeMo’s Parameter-Efficient Fine-Tuning (PEFT) support for LoRA/QLoRA techniques han2024parameter, combined with sharded checkpoints and distributed optimizers, further simplified experimentation and scale-out operations.

### v3

@@ Abstract
Figure 1: PayPal is transforming from a payment company to a commerce platform with, our commerce Paypal Agent, which leverages NVIDIA’s Nemo.
@@ 1 Introduction
The emergence of agentic artificial intelligence represents a paradigm shift in how consumers interact with e-commerce platforms Purcarea (2025); Kodali (2025), promising to transform the traditional shopping experience into an intelligent, conversational journey. PayPal’s Commerce Agent embodies this vision by enabling autonomous shopping assistance across the entire commerce lifecycle, from pre-purchase discovery through post-purchase support. This multi-agent system leverages advanced language models to provide personalized shopping experiences, positioning itself at the forefront of the agentic commerce revolution.
The complexity of modern e-commerce interactions demands sophisticated AI systems capable of understanding nuanced customer intent, navigating vast product catalogs, and orchestrating seamless transactions. Traditional rule-based systems and simple chatbots fail to meet these requirements, necessitating the development of intelligent agents powered by large language models (LLMs) Gold . However, the deployment of such systems in production environments presents significant challenges in terms of latency, cost, and performance optimization, particularly when handling the scale and diversity of PayPal’s global commerce ecosystem.
@@
We collaborated with NVIDIA to solve these performance challenges by leveraging the NeMo Framework NVIDIA Corporation (2025) and Nemotron model family with specialized, fine-tuned small language models (SLMs) optimized for commerce search tasks.
@@ 2.1 E-Commerce Information Retrieval
Traditional e-commerce information retrieval systems Manning et al. (2008) have predominantly relied on structured data representations, operating over well-defined product catalogs, taxonomies, and attribute-based schemas to enable search and recommendation functionalities. The conventional approach involves a systematic pipeline of converting unstructured textual data—such as product descriptions, user reviews, and query inputs—into structured formats through information extraction techniques, followed by conducting search operations over these normalized data structures. This methodology leverages relational databases, inverted indices, and rule-based matching algorithms to achieve precise retrieval based on explicit attribute matching and categorical filtering. While this structured approach provides deterministic results and enables efficient query processing through established indexing mechanisms, it faces inherent limitations in handling semantic nuances, contextual understanding, and the growing complexity of natural language queries that characterize modern e-commerce interactions Wang and Na (2024).
Recent advances in large language models (LLMs) have demonstrated strong capabilities in natural language understanding and generation, showing new opportunities for bridging the gap between unstructured user queries and structured product representations Gao et al. (2023). These models can effectively extract user intent and product attributes from conversational queries, and reformulate ambiguous or underspecified searches into precise, structured formats that align with product catalog schemas, thereby enhancing retrieval effectiveness in e-commerce systems Zhang et al. (2025); Jagerman et al. (2023).
@@ 2.2 LLM Agent in Search and Recommendation
LLM agents Huang et al. (2024) are transforming commerce search and recommendation systems by enabling more natural, context-aware interactions between users and e-commerce platforms. Unlike traditional keyword-based search systems that rely on exact matches and predefined filters, LLM agents leverage large language models to understand user intent across multiple dimensions, including implicit preferences, contextual constraints, and evolving needs throughout a shopping session. These agents perform sophisticated query understanding by extracting attributes (e.g., brand, color, price range, size), expanding queries with related terms and synonyms, and reformulating ambiguous requests into structured search parameters. Beyond search, LLM agents orchestrate multi-step recommendation workflows by reasoning over product catalogs, user purchase history, and real-time session data to generate personalized suggestions. They handle complex conversational scenarios such as comparison requests (”show me Nike shoes similar to what I bought last month but in blue”), constraint satisfaction (”find running shoes under $\mathdollar 100$ with good arch support”), and cross-category exploration (”I need shoes for hiking, what else should I get for my trip?”). By integrating retrieval systems with ranking models, LLM agents ground their responses in actual product inventory while maintaining conversational fluency, ultimately reducing search friction, increasing discovery, and improving conversion rates in modern commerce platforms.
@@ 2.3 NVIDIA’s NeMo Framework
NVIDIA NeMo Framework NVIDIA Corporation (2025) is a complete AI platform that helps researchers and developers work with Large Language Models (LLMs), multimodal systems, and speech AI applications. The framework supports the entire process of building AI models, from preparing data and training models to putting them into production, and can run on different computing environments including local servers, data centers, or cloud services. NeMo includes tools for data preparation, efficient training methods like supervised fine-tuning and LoRA (a technique that reduces training time and resources), and multiple options for deploying trained models.
@@
Figure 3: NeMo Framework NVIDIA Corporation (2024) NVIDIA NeMo is an end-to-end platform for developing custom generative AI, anywhere. It includes tools for training, fine-tuning, retrieval-augmented generation, guardrailing, data curation, and pretrained models. It has offerings across the tech stack, from frameworks to high-level API endpoints.
@@ 3.1 LLM-Powered Commerce Search &\& Recommendation
To leverage the powerful text generation and understanding capabilities of large language models (LLMs), we adopt the Hypothetical Document Embeddings (HyDE) Gao et al. (2023) approach for product retrieval. When a user prompt is given, the LLM first extracts (attribute, value) pairs from the natural language query and reformulates it into a structured format of (product category, attributes, values). The LLM then generates hypothetical product information that captures the user’s intent and desired product characteristics. This hypothetical product description, while potentially containing idealized or imprecise details, effectively encodes the relevance patterns of what the user is seeking. The structured query and hypothetical product information are then sent to our search engine, which retrieves matching products from our product catalog by identifying similar real products in the embedding space, effectively grounding the generated hypothetical description to actual inventory while filtering out any inaccurate details through the encoder’s dense representation.
@@ 3.2 Commerce Agent System Architecture
To further our agentic commerce initiatives, we developed a comprehensive agentic system to support conversational AI experiences for users through both web and mobile platforms. Our architecture employs a Java-based backend API server, utilizing the LangChain Mavroudis (2024) agent orchestration framework to develop intelligent products that leverage multiple language models and coordinate internal knowledge systems, domain services, and existing machine learning solutions. Our system implements a hybrid cloud solution that integrates PayPal’s internal technology stack with cloud platform services to ensure scalability, reliability, and performance for production deployment. (see Figure 4)
@@ 4.2 Evaluation Methodology
We finetuned and evaluated multiple model variants with different hyperparameter configurations. Our evaluation employs LLM-as-a-Judge Gu et al. (2024), using pairwise and rubric-based comparative evaluations for qualitative analysis, to assess model performance.
@@ 4.3 Supervised Fine Tuning (SFT)
We leveraged NVIDIA’s NeMo SDK as the foundation for fine-tuning a family of small language models (SLMs), focusing on the llama3.1-nemotron-nano-8B-v1 architecture. We conducted comprehensive experiments training 20 LoRA-based model variants while systematically sweeping hyperparameter spaces including learning rates, optimizers (Adam, AdamW), cosine annealing schedules, and LoRA ranks. The NeMo Framework SDK significantly enhanced developer productivity by providing turnkey multi-GPU and multi-dimensional parallelism capabilities, reducing time-to-train through well-maintained training recipes. NeMo’s Parameter-Efficient Fine-Tuning (PEFT) support for LoRA/QLoRA techniques Han et al. (2024), combined with sharded checkpoints and distributed optimizers, further simplified experimentation and scale-out operations.